# Incorporating Arbitrary Matrix Group Equivariance into KANs

### v2

@@ 7 Conclusion
Supplementary Material
@@ Appendix A Common matrix groups and their generators
Groups $\mathrm{SO}(2)$ and $\mathrm{O}(2)$.
@@
It corresponds to an infinite generator:
@@
Groups $\mathrm{SO}^{+}(1,3)$, $\mathrm{SO}(1,3)$, and $\mathrm{O}(1,3)$.
@@
$\begin{cases}A_{1}=\begin{bmatrix}0&1&0&0\\ 1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix},A_{2}=\begin{bmatrix}0&0&1&0\\ 0&0&0&0\\ 1&0&0&0\\ 0&0&0&0\end{bmatrix},\\ A_{3}=\begin{bmatrix}0&0&0&1\\ 0&0&0&0\\ 0&0&0&0\\ 1&0&0&0\end{bmatrix},A_{4}=\begin{bmatrix}0&0&0&0\\ 0&0&1&0\\ 0&-1&0&0\\ 0&0&0&0\end{bmatrix},\\ A_{5}=\begin{bmatrix}0&0&0&0\\ 0&0&0&1\\ 0&0&0&0\\ 0&-1&0&0\end{bmatrix},A_{6}=\begin{bmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\\ 0&0&-1&0\end{bmatrix},\end{cases}$ (27)
@@
Appendix B Proof of Theorem 1
See 1
@@ Appendix B Proof of Theorem 1
Using the structure of $U_{m}$ shown in Eq. 15 and applying the rules from Eq. 4, we can derive the group representation of $U_{m}$:
@@
where $\rho_{i}$ is the group representation of $U_{i}$. Combining Eqs. 30 and 31, we have:
@@
Note that the group transformation in the scalar space $T_{0}$ is the identity transformation, then we can obtain the group representation of $U_{gi}$ from Eq. 11:
@@
where $\rho_{i,a}$ is the group representation of $T(p_{i,a},q_{i,a})$. Therefore, applying the group transformation to $v_{gi}$ in Eq. 12 results in:
@@
Substitute Eq. 34 into Eq. 14:
$f_{b}(\rho_{gi}(g)v_{gi})=$
$\begin{cases}\left[\bigoplus_{a=1}^{c_{i}}s_{i,a}B_{b}(s_{i,a})\right]\oplus\left[\bigoplus_{a=1}^{A_{i}}\rho_{i,a}(g)v_{i,a}B_{b}(s_{i,a}^{\prime})\right],\\ \hskip 170.0ptb<G+k,\\ \left[\bigoplus_{a=1}^{c_{i}}s_{i,a}\mathrm{silu}(s_{i,a})\right]\oplus\left[\bigoplus_{a=1}^{A_{i}}\rho_{i,a}(g)v_{i,a}\mathrm{silu}(s_{i,a}^{\prime})\right],\\ \hskip 170.0ptb=G+k.\end{cases}$ (35)
Similar to Eq. 33, we can derive the group representation of $U_{i}$ from Eq. 10:
@@
Note that the right-hand side of Appendix B is the result of applying $\rho_{i}(g)$ to $f_{b}(v_{gi})$, which means:
@@
Substitute Eq. 37 into Eq. 30:
@@
Combining Eqs. 32 and 38, Eq. 16 is proven. ∎
Appendix C Implementation details
C.1 Particle scattering
In particle scattering, we generate training sets of different sizes, and the corresponding test sets have the same sizes as the training sets. We train EKAN using the Adan optimizer [54] with the learning rate of $3\times 10^{-3}$ and the batch size of $500$. For datasets with the training set size $<1000$, we set the number of epochs to $7000$, while for datasets with the training set size $\geq 1000$, we set the number of epochs to $15000$, which is sufficient for the MSE loss to converge to the minimum. We perform this experiment on a single-core NVIDIA GeForce RTX 3090 GPU with available memory of $24576$ MiB. For each data point, the inference time of EKAN is approximately $18$ ms.
C.2 Three-body problem
In the three-body problem, we control the number of parameters by adjusting the middle layer width $N$ of the model. We list the correspondence between the model’s shape and the number of parameters in Tab. 4. We train all models using the Adan optimizer [54] with the learning rate of $3\times 10^{-3}$, the batch size of $500$, and for $5000$ epochs. The grids of EKAN and KAN are updated every $5$ epochs and stop updating at the $50$th epoch. We perform this experiment on a single-core NVIDIA GeForce RTX 3090 GPU with available memory of $24576$ MiB. The relationship between EKAN’s inference time per data point and the number of parameters is shown in Tab. 5.
@@ C.2 Three-body problem
Number of parameters $10^{4.5}$ $10^{4.75}$ $10^{5}$ $10^{5.25}$ $10^{5.5}$
Time (ms) $7$ $12$ $13$ $14$ $16$
Table 5: The relationship between EKAN’s inference time per data point and the number of parameters.
C.3 Top quark tagging
In top quark tagging, we train all models using the Adan optimizer [54] with the learning rate of $3\times 10^{-3}$ and the batch size of $500$. For datasets with the training set size $\leq 1000$, we set the number of epochs to $1000$, while for datasets with the training set size $>1000$, we set the number of epochs to $2000$, which is sufficient for the BCE loss to converge to the minimum. The grids of EKAN and KAN are updated every $5$ epochs and stop updating at the $50$th epoch. We perform this experiment on a single-core NVIDIA GeForce RTX 3090 GPU with available memory of $24576$ MiB. For each data point, the inference time of EKAN is approximately $13$ ms.

### v3

@@ 7 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Common matrix groups and their generators
A.1 Groups $\mathrm{SO}(2)$ and $\mathrm{O}(2)$
@@
It corresponds to an infinitesimal generator:
@@
A.2 Groups $\mathrm{SO}^{+}(1,3)$, $\mathrm{SO}(1,3)$, and $\mathrm{O}(1,3)$
@@
$\begin{cases}A_{1}=\begin{bmatrix}0&1&0&0\\ 1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix},A_{2}=\begin{bmatrix}0&0&1&0\\ 0&0&0&0\\ 1&0&0&0\\ 0&0&0&0\end{bmatrix},A_{3}=\begin{bmatrix}0&0&0&1\\ 0&0&0&0\\ 0&0&0&0\\ 1&0&0&0\end{bmatrix},\\ A_{4}=\begin{bmatrix}0&0&0&0\\ 0&0&1&0\\ 0&-1&0&0\\ 0&0&0&0\end{bmatrix},A_{5}=\begin{bmatrix}0&0&0&0\\ 0&0&0&1\\ 0&0&0&0\\ 0&-1&0&0\end{bmatrix},A_{6}=\begin{bmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\\ 0&0&-1&0\end{bmatrix},\end{cases}$ (24)
@@
Appendix B Equivariant Multi-Layer Perceptrons (EMLP)
EMLP (Finzi et al., 2021) embeds matrix group equivariance into MLPs layerwise. Given the input space $U_{i}$ and output space $U_{o}$, the linear weight matrix $W\in U_{o}\otimes U_{i}^{*}$ should satisfy Equation 5, i.e., $\forall g\in\widetilde{G},v_{i}\in U_{i}:\rho_{o}(g)Wv_{i}=W\rho_{i}(g)v_{i}$. So the coefficients of each term in $v_{i}$ are equal $\forall g\in\widetilde{G}:\rho_{o}(g)W=W\rho_{i}(g)$. Flattening the linear weight matrix $W$ into a vector, we have $\forall g\in\widetilde{G}:\left[\rho_{o}(g)\otimes\rho_{i}(g^{-1})^{\top}\right]\mathrm{vec}(W)=\mathrm{vec}(W)$. Combined with Equation 4, the linear weight matrix $W$ is invariant in the space $U_{o}\otimes U_{i}^{*}$:
where $\rho_{o,i}=\rho_{o}\otimes\rho_{i}^{*}$ is the group representation of $U_{o}\otimes U_{i}^{*}$. Decomposing the group representation $\rho_{o,i}(g)$ into discrete and infinitesimal generators as shown in Equation 2, Equation 27 is equivalent to the following constraint:
Appendix C Proof of Theorem 4.1
See 4.1
@@ Appendix B Equivariant Multi-Layer Perceptrons (EMLP)
Using the structure of $U_{m}$ shown in Equation 13 and applying the rules from Equation 4, we can derive the group representation of $U_{m}$:
@@
where $\rho_{i}$ is the group representation of $U_{i}$. Combining Equations 29 and 30, we have:
@@
Note that the group transformation in the scalar space $T_{0}$ is the identity transformation, then we can obtain the group representation of $U_{gi}$ from Equation 9:
@@
where $\rho_{i,a}$ is the group representation of $T(p_{i,a},q_{i,a})$. Therefore, applying the group transformation to $v_{gi}$ in Equation 10 results in:
@@
Substitute Equation 33 into Equation 12:
$f_{b}(\rho_{gi}(g)v_{gi})=\begin{cases}\left[\bigoplus_{a=1}^{c_{i}}s_{i,a}B_{b}(s_{i,a})\right]\oplus\left[\bigoplus_{a=1}^{A_{i}}\rho_{i,a}(g)v_{i,a}B_{b}(s_{i,a}^{\prime})\right],\quad b<G+k,\\ \left[\bigoplus_{a=1}^{c_{i}}s_{i,a}\mathrm{silu}(s_{i,a})\right]\oplus\left[\bigoplus_{a=1}^{A_{i}}\rho_{i,a}(g)v_{i,a}\mathrm{silu}(s_{i,a}^{\prime})\right],\quad b=G+k.\end{cases}$ (34)
Similar to Equation 32, we can derive the group representation of $U_{i}$ from Equation 8:
@@
Note that the right-hand side of Equation 34 is the result of applying $\rho_{i}(g)$ to $f_{b}(v_{gi})$, which means:
@@
Substitute Equation 36 into Equation 29:
@@
Combining Equations 31 and 37, Equation 14 is proven. ∎
Appendix D Grid update
Similar to KANs (Liu et al., 2024b), EKAN updates grids based on the input activations. At the same time, the linear weights should also be updated in order to keep the output features unchanged. Let the post-activation values of the grids before and after the update be denoted as $V_{m},V_{m}^{\prime}\in\mathbb{R}^{N\times(G+k+1)d_{i}}$, where $N$ is the number of samples. We first project the linear weight blocks into the equivariant subspace as $\mathrm{vec}(\widetilde{W}_{b})=P\mathrm{vec}(W_{b})$ and compute the output activations $V_{go}=V_{m}\widetilde{W}^{\top}=\sum_{b=0}^{G+k}V_{m,b}\widetilde{W}_{b}^{\top}$, where $P$ is the equivariant projector obtained from Section 4.3. Then, the updated equivariant linear weights $\widetilde{W}^{\prime}$ should satisfy $V_{go}=V_{m}^{\prime}\widetilde{W}^{\prime\top}$, and we have $\widetilde{W}^{\prime}=V_{go}^{\top}(V_{m}^{\prime\top})^{\dagger}$. We finally restore the updated linear weight blocks $\mathrm{vec}(W_{b}^{\prime})=P^{\dagger}\mathrm{vec}(\widetilde{W}_{b}^{\prime})$.
Appendix E Implementation details
E.1 Particle scattering
In particle scattering, we generate training sets of different sizes, and the corresponding test sets have the same sizes as the training sets. We train EKAN using the Adan optimizer (Xie et al., 2024) with the learning rate of $3\times 10^{-3}$ and the batch size of $500$. For datasets with the training set size $<1000$, we set the number of epochs to $7000$, while for datasets with the training set size $\geq 1000$, we set the number of epochs to $15000$, which is sufficient for the MSE loss to converge to the minimum. We perform this experiment on a single-core NVIDIA GeForce RTX 3090 GPU with available memory of $24576$ MiB.
E.2 Three-body problem
In the three-body problem, we control the number of parameters by adjusting the middle layer width $N$ of the model. We list the correspondence between the model’s shape and the number of parameters in Table 4. We train all models using the Adan optimizer (Xie et al., 2024) with the learning rate of $3\times 10^{-3}$, the batch size of $500$, and for $5000$ epochs. The grids of EKAN and KAN are updated every $5$ epochs and stop updating at the $50$th epoch. We perform this experiment on a single-core NVIDIA GeForce RTX 3090 GPU with available memory of $24576$ MiB.
@@ E.2 Three-body problem
E.3 Top quark tagging
In top quark tagging, we train all models using the Adan optimizer (Xie et al., 2024) with the learning rate of $3\times 10^{-3}$ and the batch size of $500$. For datasets with the training set size $\leq 1000$, we set the number of epochs to $1000$, while for datasets with the training set size $>1000$, we set the number of epochs to $2000$, which is sufficient for the BCE loss to converge to the minimum. The grids of EKAN and KAN are updated every $5$ epochs and stop updating at the $50$th epoch. We perform this experiment on a single-core NVIDIA GeForce RTX 3090 GPU with available memory of $24576$ MiB.